OpenStudy (anonymous):

what is 499800 expressed to 2 significant figures? Would it be 500000? or 490000? And what would that be expressed in KiloPascals?

OpenStudy (aum):

I am thinking \(\large 5.0 \times 10^5\) (2 significant figures)

OpenStudy (anonymous):

Thank you Mayan but the question is asking me to express it in 2 siginificant figures. And aum so, I WOULD round up then which would give 500,000 but if that's the case that's not 2 significant figures. So Do you think the question just wants it to be shown with ATLEAST 2 significant figures?

OpenStudy (anonymous):

Thank you both, I would give both of you best answer but I can't, do one of you want it more than the other?

OpenStudy (aum):

If you round 499800 to 500000 it has only one significant figure (5). But if you express it in scientific notation with one decimal place it will be two significant figures: \(\large 5.0 \times 10^5\) (2 significant figures)

OpenStudy (anonymous):

hmmm okay I see what you're saying, I didn't think the scientific notation portion counted as being a significant figure, I thought it just served as a place holder, but okay thank you!

OpenStudy (aum):

\(\large 499800 = 4.998 \times 10^5\) (4 significant figures). Round it to two significant figures: \(\large 5.0 \times 10^5\)

OpenStudy (anonymous):

alright thank you aum